OpenStudy (anonymous):

5x/4+1/2=0... please help me solve the problem step by step .... DONT JUST GIVE ME THE ANSWER

OpenStudy (samigupta8):

5x/4=-1/2 X=-2/5

OpenStudy (loser66):

@samigupta8 you failed!! :) @dixiegirl8794 subtract 1/2 both sides, what do you get?

OpenStudy (anonymous):

I got the answer on my own.. thanks yall